Related Rates: Problem Solving


Direction: Solve each word problem involving a related rates. Make a sketch of your work

if necessary, and then apply differentiation.


Sand is being dropped at a rate of 10 cubic feet per minute onto a conical pile. If

the height of the pile is always twice the base radius, at what rate is the height

increasing when the pile is 8 feet high?


PLEASE ANSWER MY QUESTION QUICKLY!!

DEADLINE : 05/03/2022 11 : 00 PM




1
Expert's answer
2022-05-08T13:41:41-0400
"V=\dfrac{1}{3}\pi r^2h"

Given "h=2r"


"V=\dfrac{\pi}{12}h^3"

Differentiate both sides with respect to "t"


"\dfrac{dV}{dt}=\dfrac{\pi}{4}h^2(\dfrac{dh}{dt})"

Then


"\dfrac{dh}{dt}=\dfrac{4}{\pi h^2}(\dfrac{dV}{dt})"

When the pile is 8 feet high


"\dfrac{dh}{dt}=\dfrac{4}{\pi (8ft)^2}(10{ft}^3/min)"

"=\dfrac{5}{8\pi}ft/min\approx0.199ft/min"

The height is increasing at rate of 0.199 feet per minute when the pile is 8 feet high.